# SeamlessEdit: Background Noise Aware Zero-Shot Speech Editing with in-Context Enhancement

### v1

@@ 4 Conclusion
This study presents a novel speech editing framework that is robust to noise scenarios. It overcomes the limitations of conventional methods confined to controlled acoustic conditions. By handling diverse background noises and preserving ambient audio, it significantly expands the applicability in noisy recordings. The proposed framework not only ensures the intelligibility of the edited speech signal but also maintains its naturalness and fidelity. This advancement makes speech editing even more feasible for real-world applications such as podcast editing, interview restoration, and archival enhancement.

### v2

@@ 4 Conclusion
This work presents SeamlessEdit, a noise-robust speech editing framework designed for real-world conditions. Unlike prior methods limited to clean studio recordings, SeamlessEdit can handle diverse background noises while preserving useful ambience, ensuring intelligibility and naturalness of edited speech. By jointly modeling content, prosody, and environmental acoustics, it enables seamless word- and phrase-level editing without audible artifacts. Applications include podcast production, interview restoration, and archival enhancement, highlighting its practicality in everyday and professional audio workflows.